# Lieb-Schultz-Mattis-Type and Laughlin-Type Argument for the Quantum Hall Effect in Lattice Fermions with Spiral Boundary Conditions

### v1

@@ Abstract
We derive the condition for the occurrence of the quantum Hall effect in two-dimensional lattice systems, expressed as $\phi\nu-\rho\in\mathbb{Z}$, where $\phi$, $\nu$, and $\rho$ denote the magnetic flux, the Chern number, and the electron density, respectively. By employing spiral boundary conditions, which treat the system as an extended one-dimensional chain, this condition is obtained directly through a Lieb-Schultz-Mattis-type and Laughlin-type argument. This approach improves upon the preceding work based on conventional periodic boundary conditions, where the condition was derived indirectly containing redundant system-size dependence. The key to this approach is that the spatial directions of the external force and the response can be manipulated by a factor of the system size.
@@
Lu, Ran and Oshikawa derived the conditions for the appearance of the QHE in lattice systems with interactions by combining the LSM-type argument with Laughlin-type argument as[23, 24]
@@
where $\phi$, $\nu$, and $\rho$ denote the magnetic flux, the Chern number, and the electron density, respectively. In their theory based on periodic boundary conditions (PBCs), the condition for QHE appears indirectly passing through the intermediate relation which depends on an artificial geometric parameter. Therefore, improvement of this proof is desirable to eliminate this redundancy.
@@
Figure 1: Several boundary conditions for 2D square lattices with $L_{1}\times L_{2}$ sites: (a) Conventional periodic boundary conditions (PBCs), (b) Spiral boundary conditions (SBCs). For the systems with SBCs, the lattices are labeled as extended 1D chains. $\bm{A}_{i}$ and $\bm{A}_{i}^{\prime}$ are primitive vectors.
@@
where $c^{\mathstrut}_{m,n}$ ($c_{m,n}^{{\dagger}}$) is the annihilation (creation) operator at $(m,n)$ site, $t$ is the hopping amplitude, $V(\{n_{m,n}\})$ is the fermion interactions consists of the number operator $n_{m,n}\equiv c_{m,n}^{{\dagger}}c_{m,n}^{\mathstrut}$. The gauge field is introduced in the Landau gauge as
@@
which give the magnetic flux $B=-\frac{2\pi\hbar}{ea^{2}}\phi$ with
@@
Figure 2: The gauge field $\theta^{\alpha}$ in a square lattice in a magnetic field with (a) periodic boundary conditions, and (b) spiral boundary conditions.
In the absence of a magnetic field ($\phi=0$), the Hamiltonian possesses translational symmetry by one lattice spacing along both directions, $\mathcal{H}=T_{i}^{\mathstrut}\mathcal{H}T_{i}^{-1}$. For $\phi\neq 0$, this translational symmetry is broken due to the position dependence of $\theta^{2}_{m,n}$ and sublattice structure appears as shown in Fig. 3. Therefore, $L_{1}$ is selected so that $L_{1}/q\in\mathbb{Z}$. On the other hand, a similar symmetry is recovered as $\mathcal{H}=\tilde{T}_{i}^{\mathstrut}\mathcal{H}\tilde{T}_{i}^{-1}$ by introducing the magnetic translation operators
@@
Figure 3: Sublattice structure of the system $(L_{1},L_{2},q)=(6,4,3)$ with a magnetic field ($\phi=p/q$). Here, $L_{1}/q$ is chosen to be an integer.
@@
where $\mathcal{T}$ denotes time ordering, $T$ is a sufficiently long period, and $\mathcal{H}_{i}(\Phi)\equiv U_{i}^{\Phi/2\pi}\mathcal{H}U_{i}^{-\Phi/2\pi}$.
@@
This is also an expression of the Laughlin’s argument for QHE. Since $F_{i}(2\pi)$ plays a similar role to $U_{i}$, we can introduce a magnetic translation operator as an analogy of Eq. (6),
@@
Thus, following an LSM-type argument, the condition for $z(\Phi)\neq 0$ becomes
@@
Figure 4: Fluxes $\Phi_{i}$ for 2D lattice applied as $\mathcal{H}_{i}(\Phi)=U_{i}^{\Phi/2\pi}\mathcal{H}U_{i}^{-\Phi/2\pi}$. In the system with SBCs, the insertion of flux $\mathcal{H}_{1}^{\prime}(\Phi_{1})$ has the same effect of the flux $\mathcal{H}_{2}^{\prime}(L_{2}\Phi_{1})$ due to the relation $U_{1}^{\prime}=U_{2}^{\prime L_{2}}$.
@@
For $i=2$, the exponent part of the polarization operator is confirmed as that of SBCs as
@@
$T_{1}^{\prime}U_{2}^{\prime}T_{1}^{\prime-1}=T_{1}^{\prime}U_{1}^{\prime 1/L_{2}}T_{1}^{\prime-1}={\rm e}^{-{\rm i}2\pi\hat{\rho}}U_{2}^{\prime}.$ (34)
We choose $U_{2}^{\prime}$ rather than $U_{1}^{\prime}$ here because it eliminates the $L_{2}$-dependence from this relation. This was introduced for an extention of the LSM theorem to higher dimensions[16]. The magnetic translation operators are
@@
and the Laughlin-type relation becomes
@@
Since, in the system with SBCs, the insertion of flux $\mathcal{H}_{1}^{\prime}(\Phi_{1})$ has the same effect of the flux $\mathcal{H}_{2}^{\prime}(L_{2}\Phi_{1})$ due to the relation of Eq. (33), the magnetic translation operator corresponding to Eq. (17) is
@@
In this work, we have derived the condition for the appearance of the quantum Hall effect (QHE) in 2D lattice systems $\phi\nu-\rho\in\mathbb{Z}$, by means of a Lieb-Schultz-Mattis (LSM)-type and Laughlin-type argument. By introducing the spiral boundary conditions (SBCs), we have successfully eliminated the artificial dependence on the system width $L_{2}$ that arises in the conventional argument based on periodic boundary conditions (PBCs). Under the SBC, the 2D lattice can be mapped onto an extended 1D chain, so that explicit dependence on the cross section can be avoided, and the effects of flux insertion in different directions can be treated on equal footing. This construction provides a more transparent physical picture of the quantization of the Hall conductance in lattice fermion systems.
The key to this approach is that the spatial directions of the external force and the response can be manipulated by a factor of the system size as in Eqs. (33)-(35), and (38), thereby allowing the flux insertion process to be represented without introducing any redundant parameters. Consequently, the condition $\phi\nu-\rho\in\mathbb{Z}$ emerges directly from the magnetic translation symmetry, demonstrating that the quantization of the Hall conductance can be understood purely from symmetry and topology, without relying on explicit band-structure calculations.
Our approach unifies the LSM argument for interacting systems and Laughlin’s gauge argument for the QHE, and provides a useful framework for analyzing correlated topological phases in lattice models. Since the present formalism is formulated entirely in terms of the many-body wave function and the polarization operators, it can be straightforwardly extended to systems with interactions, disorder, or nontrivial lattice geometries. Applications to Chern insulators, fractional Chern insulators, and topological superconductors under SBCs will be interesting future problems.

### v2

@@ Abstract
We derive the condition for the occurrence of the integer quantum Hall effect in two-dimensional lattice systems with interactions, expressed as $\phi\nu-\rho\in\mathbb{Z}$, where $\phi$, $\nu$, and $\rho$ denote the magnetic flux, the Chern number, and the electron density, respectively. By employing spiral boundary conditions, which treat the system as an extended one-dimensional chain, this condition is obtained directly through a Lieb-Schultz-Mattis-type and Laughlin-type argument. This approach improves upon the preceding work based on conventional periodic boundary conditions, where the condition was derived indirectly with redundant system-size dependence. The key to this approach is that the spatial directions of the external force and the response can be systematically controlled by a factor of the system size.
@@
Lu, Ran, and Oshikawa derived the condition for the appearance of the QHE in lattice systems with interactions by combining the LSM-type argument with Laughlin-type argument as[23, 24]
@@
where $\phi$, $\nu$, and $\rho$ denote the magnetic flux, the Chern number, and the electron density, respectively. In their theory based on periodic boundary conditions (PBCs), the condition for QHE appears indirectly passing through the intermediate relation which depends on an artificial geometric parameter. Therefore, it is desirable to improve the proof to eliminate this redundancy.
@@
Figure 1: (Color) Boundary conditions for 2D lattices: (a) Conventional periodic boundary conditions (PBCs), (b) Spiral boundary conditions (SBCs), where the system is described as an extended 1D chain. Fluxes $\Phi_{i}$ for 2D lattices are applied as $\mathcal{H}_{i}(\Phi)=U_{i}^{\Phi/2\pi}\mathcal{H}U_{i}^{-\Phi/2\pi}$. For SBCs, the insertion of flux $\mathcal{H}_{1}^{\prime}(\Phi_{1})$ has the same effect as the flux $\mathcal{H}_{2}^{\prime}(L_{2}\Phi_{1})$ due to the relation $U_{1}^{\prime}=U_{2}^{\prime L_{2}}$.
@@
where $c^{\mathstrut}_{m,n}$ ($c_{m,n}^{{\dagger}}$) is the annihilation (creation) operator at $(m,n)$ site, $t$ is the hopping amplitude, $V(\{n_{m,n}\})$ represents the fermion interaction and is written in terms of the number operator $n_{m,n}\equiv c_{m,n}^{{\dagger}}c_{m,n}^{\mathstrut}$. The gauge field is introduced in the Landau gauge as
@@
which gives the magnetic flux $B=-\frac{2\pi\hbar}{ea^{2}}\phi$ with
@@
In the absence of a magnetic field ($\phi=0$), the Hamiltonian possesses translational symmetry by one lattice spacing along both directions, $\mathcal{H}=T_{i}^{\mathstrut}\mathcal{H}T_{i}^{-1}$. For $\phi\neq 0$, this translational symmetry is broken due to the position dependence of $\theta^{2}_{m,n}$ and sublattice structure appears as shown in Fig. 3. Therefore, $L_{1}$ is selected so that $L_{1}/q\in\mathbb{Z}$. On the other hand, a similar symmetry is satisfied as $\mathcal{H}=\tilde{T}_{i}^{\mathstrut}\mathcal{H}\tilde{T}_{i}^{-1}$ by introducing the magnetic translation operators
@@
Figure 2: (Color) The gauge field $\theta^{\alpha}$ in a square lattice in a magnetic field with (a) periodic boundary conditions, and (b) spiral boundary conditions.
Figure 3: (Color) Sublattice structure of the system $(L_{1},L_{2},q)=(6,4,3)$ with a magnetic field ($\phi=p/q$). Here, $L_{1}/q$ is chosen to be an integer.
@@
This is also an expression of Laughlin’s argument for QHE [4] in lattice systems. It states that inserting a unit flux $\Phi=2\pi h/e$ in the $x_{2}$ direction pumps an integer number of fermions in the $x_{1}$ direction when $\sigma_{xy}$ is quantized.
@@
where $\mathcal{T}$ denotes time ordering, $T$ is a sufficiently long period compared with the inverse of the energy gap, and
$\mathcal{H}_{i}(\Phi)\equiv U_{i}^{\Phi/2\pi}\mathcal{H}U_{i}^{-\Phi/2\pi}.$ (16)
@@
Since $F_{i}(2\pi)$ plays a similar role to $U_{i}$, we can introduce a magnetic translation operator as an analogy of Eq. (6),
@@
Thus, following an LSM-type argument for the symmetry operation, the condition for $z(\Phi)\neq 0$ becomes
@@
For $i=2$, the exponent part of the polarization operator can be written in the SBC representation as
@@
Figure 4: (Color) Boundary conditions for 2D square lattices with $L_{1}\times L_{2}$ sites: (a) Periodic boundary conditions (PBCs), (b) Spiral boundary conditions (SBCs). For the systems with SBCs, the lattices are labeled as extended 1D chains. $\bm{A}_{i}$ and $\bm{A}_{i}^{\prime}$ are primitive vectors. When we cut the system (by green lines), the difference in the number of removed bonds in both boundary conditions is only one.
@@
$T_{1}^{\prime}U_{2}^{\prime}T_{1}^{\prime-1}={\rm e}^{-{\rm i}2\pi\hat{\rho}}U_{2}^{\prime}.$ (35)
We choose $U_{2}^{\prime}$ rather than $U_{1}^{\prime}$ here because it eliminates the $L_{2}$-dependence from this relation. This was introduced for an extension of the LSM theorem to higher dimensions[16]. The magnetic translation operators are
@@
and the Laughlin-type relation (17) becomes
@@
In Eq. (37), the Hamiltonian with a flux $\mathcal{H}_{i}^{\prime}(\Phi_{i})$ is described similarly to the case of PBC, however, the definition Eq. (16) for SBCs means that effects of the fluxes $\Phi_{1}$ and $\Phi_{2}$ are no longer independent (see also Fig. 1(b)), because $U_{1}^{\prime}$ and $U_{2}^{\prime}$ have the same phase variable except for their coefficients as in Eq. (33). Therefore, the insertion of flux $\mathcal{H}_{1}^{\prime}(\Phi_{1})$ has the same effect as the flux $\mathcal{H}_{2}^{\prime}(L_{2}\Phi_{1})$ due to the relation of Eq. (34), so that the magnetic translation operator corresponding to Eq. (18) is
@@
In this work, we have derived the condition for the appearance of the integer quantum Hall effect (QHE) in 2D lattice systems with arbitrary interactions, namely $\phi\nu-\rho\in\mathbb{Z}$, by means of a Lieb-Schultz-Mattis (LSM)-type and Laughlin-type argument. By introducing the spiral boundary conditions (SBCs), we have successfully eliminated the artificial dependence on the system width $L_{2}$ that arises in the conventional argument based on periodic boundary conditions (PBCs). Under the SBC, the 2D lattice can be mapped onto an extended 1D chain, so that explicit dependence on the system’s cross section is avoided, and the effects of flux insertion in different directions are treated on equal footing. This construction provides a more transparent physical picture of the quantization of the Hall conductance in lattice fermion systems.
The key to this approach is that the spatial directions of the external force and the response can be controlled by a factor of the system size as in Eqs. (34)-(36), and (39), thereby allowing the flux insertion process to be represented without introducing any redundant parameters. Consequently, the condition $\phi\nu-\rho\in\mathbb{Z}$ emerges directly from the magnetic translation symmetry, demonstrating that the quantization of the Hall conductance can be understood purely from symmetry and topology, without relying on explicit band-structure calculations.
Although we emphasize the advantages of the 1D description by SBCs, the system remains physically equivalent to the 2D lattice, and hoppings in the $x_{2}$ direction still appear as long-range hopping terms $c^{\dagger}_{j+L_{1}}{\rm e}^{{\rm i}\theta_{j}^{2}}c^{\mathstrut}_{j}+\mbox{H.c.}$ in the 1D chain representation. Thus, cutting the torus under SBCs removes not only the bond on the 1D chain, but also $L_{1}$ long-range bonds as shown in Fig. 4. Therefore, compared with PBCs, the number of removed bonds differs only by one, which does not affect the physics in the thermodynamic limit. Hence, the edge-state properties under the two boundary conditions are essentially the same, and Laughlin’s relation holds in both cases.
Our approach unifies the LSM argument for interacting systems and Laughlin’s gauge argument for the QHE, and provides a useful framework for analyzing correlated topological phases in lattice models. Since the present formalism is formulated entirely in terms of the many-body wave function and the polarization operators, it can be straightforwardly extended to systems with disorder or nontrivial lattice geometries, and also to the fractional QHE[23]. In addition to the present method for the QHE condition, an alternative formulation has also been proposed by Matsugatani et al.[24] in terms of many-body Chern number based on Berry connections [1, 29] and point-group symmetries. It would be also interesting to reformulate their framework using SBCs.